A researcher claims that 20 year old women on a special diet will have an average weight of 110 pounds. A sample of 15 women has an average weight of 112.5 pounds and a standard deviation of 5 pounds. A α = .01, can the claim be rejected?


1
Expert's answer
2021-12-14T06:10:50-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=110"

"H_1:\mu\not=110"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01,"

"df=n-1=15-1=14" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.976842."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.976842\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}==\dfrac{112.5-110}{5/\sqrt{15}}\approx1.9365"

Since it is observed that "|t| = 1.9365< 2.976842= t_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed "df=14" degrees of freedom, "t=1.9365" is "p =0.073257," and since "p = 0.073257 \ge 0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than "110," at the "\alpha = 0.01" significance level.